OpenStudy (anonymous):

Classify the discontinuities (if any) for the given function:f(x)={(x+4)/(x^2-3x-28) x not equal 4, x not equal 7 {-1/11 otherwise

OpenStudy (anonymous):

same as \[\frac{1}{x-7}\]

OpenStudy (anonymous):

the discontinuity at 7 is infinite, i.e.it has a vertical asymptote there

OpenStudy (anonymous):

why not removable at x=-4?

OpenStudy (anonymous):

the discontinuity at-4 is evidently removable, because i just removed it.

OpenStudy (anonymous):

haha, let me try again :-)

OpenStudy (anonymous):

it is removable at -4, but not at 7

OpenStudy (anonymous):

can you pls tell me how it is removable at -4?

OpenStudy (anonymous):

we removed it when replacing \[\frac{(x+4)}{(x^2-3x-28)}\] by \[\frac{1}{x-7}\]

OpenStudy (anonymous):

I got (x+4)/((x-7)(x+4))

OpenStudy (anonymous):

doesn't that imply a hole at x=-4?

OpenStudy (anonymous):

more precisely, your function was \[\frac{x+4}{x^2-3x-28}=\frac{1}{x-7}\]

OpenStudy (anonymous):

yes it has a hole at -4

OpenStudy (anonymous):

but if you define it to be \[-\frac{1}{11}\] the hole is filled up. that is why it is called "removable" because you have just removed it

OpenStudy (anonymous):

oh!!! I never knew that! thank you so much!

OpenStudy (anonymous):

it doesn;t matter that it says 'otherwise' by -1/11?

OpenStudy (anonymous):

yw. math often names thinks what they are. not always of course. but in this case "removable' means you can remove it, "jump" means there is a jump, and "infinite" means it goes to infinity

OpenStudy (anonymous):

ahh "otherwise"

OpenStudy (anonymous):

interesting

OpenStudy (anonymous):

otherwise = non-integers (fractions, decimals, irrational numbers,...)

OpenStudy (anonymous):

that means you have defined \[f(x)=\frac{x+4}{x^2-3x-28}\] unless \[x=-4,7\] and \[f(-4)=-\frac{1}{11}\] also \[f(7)=-\frac{1}{11}\]

OpenStudy (anonymous):

now \[f(-4)=-\frac{1}{11}\] fills the whole so removes the discontinuity there

OpenStudy (anonymous):

however \[f(7)=-\frac{1}{11}\] doesn't fill a hole because there is no hole there. function goes to infinity. i just defines the function there

OpenStudy (anonymous):

how is f(7) = -1/11? when I plug in I do not get that?

OpenStudy (anonymous):

wait

OpenStudy (anonymous):

you can define a function any way you like

OpenStudy (anonymous):

understand the filling of a hole now!

OpenStudy (anonymous):

i can say \[f(x)=x^2\] unless x = 2, and \[f(2)=50\]

OpenStudy (anonymous):

good!

OpenStudy (anonymous):

filling in the hole means just that. plugging it up

OpenStudy (anonymous):

so if the limit of my function is \[-\frac{1}{11}\] at x = -4 i can define it to be \[-\frac{1}{11}\] at -4 and that fills the hole

OpenStudy (anonymous):

but don't forget i can define a function any way i like. i can say \[f(x)=\frac{1}{x}\] and \[f(0)=5\]

OpenStudy (anonymous):

the fact that the formula \[\frac{1}{x}\] has no meaning at x = 0 doesn't stop me from defining my function to be whatever i like at x = 0

OpenStudy (anonymous):

so for f(7) - -1/11 you are not plugging in but rather defining

OpenStudy (anonymous):

exactly!

OpenStudy (anonymous):

*so for f(7) =-1/11 you are not plugging in but rather defining

OpenStudy (anonymous):

right. just saying it.

OpenStudy (anonymous):

that is rather awesome!

OpenStudy (anonymous):

thanks!!!

OpenStudy (anonymous):

that is one thing great about math. it is my function, i can define it as i choose

OpenStudy (anonymous):

yw